# Keeping Medical AI Healthy: A Review of Detection and Correction Methods for System Degradation

### v1

@@ Abstract
AI in Healthcare, AI Performance Degradation, AI Self-Correction, Data Drift, Model Drift, Clinical Decision Support, AI Monitoring, AI Reliability, AI Safety
@@ 1 Introduction
The world is seeing increasing applications of Artificial Intelligence (AI) in healthcare [1], transforming disease diagnosis, patient monitoring, outcome prediction, and treatment planning. AI technologies range from image analysis algorithms in radiology [2] to predictive models that leverage electronic health records (EHR) for early disease detection [3]. By March 2025, the United States Food and Drug Administration (FDA) has listed 1,016 AI-enabled medical devices that have met the FDA’s applicable premarket requirements [4], and the number of AI/ML-based algorithms approved for clinical use by FDA continues to grow at a rapid rate [5]. As these AI systems continue to be adopted, ensuring their reliability and sustained performance becomes critically important. A recent study [6] analyzed 100 commercially available AI products in radiology and found that while AI adoption is accelerating, many systems lack rigorous scientific validation. Specifically, 64% of the products had no peer-reviewed evidence, and only a small fraction demonstrated clinical impact beyond diagnostic accuracy. The reliability of AI models in healthcare is not merely a technological concern but a crucial factor in ensuring patient safety and quality of care. Inaccurate or degraded AI performance can lead to severe consequences, including misdiagnosis, inappropriate treatment recommendations, and a loss of trust from healthcare providers and patients.
Figure 1: Illustration of performance degradation of medical AI models due to data or model shifts.
@@
Another type of variation is within-environment variation, also known as temporal variation. In this scenario, an AI model is designed and deployed within a fixed environment, meaning the external setting remains constant. However, over time, the model still experiences performance degradation. This phenomenon can be vividly depicted as “AI aging” [10], where the model’s effectiveness declines due to shifts in the underlying data distribution. Such variations can occur abruptly, triggering a sharp drop in performance that can be classified as an anomaly event. Conversely, a more subtle but persistent issue arises when the AI model suffers from a gradual performance decline, which may go unnoticed without continuous monitoring.
Only recently has the research community and industry started recognizing the importance of monitoring AI performance degradataion [11, 12, 13]. Chen et al. [14] demonstrate that clinical decision support systems trained on historical EHR data experience significant declines in predictive accuracy as the underlying clinical practice evolves, suggesting a natural decay in the utility of past clinical data. Their findings show that AI models relying on older datasets performed worse than those trained on more recent data, indicating an inherent susceptibility to performance degradation.
The study [15] empirically evaluates the long-term performance degradation of ML models predicting in-hospital mortality at the time of emergency admission, using a real-world dataset of 1.83 million patient discharge records from Maryland State Inpatient Database (2016–2018). By analyzing four top-performing ML models over 2.5 years, the authors track degradation across key metrics, including Area Under the Receiver Operator Characteristic Curve (AUC), accuracy, precision, and recall. The findings reveal that while ML models can remain effective for over a year post-training, gradual performance decline necessitates strategic retraining. Additionally, the study highlights the limitations of 10-fold cross-validation in predicting long-term reliability and emphasizes the need for adaptive monitoring to maintain robustness in clinical AI applications.
@@
Another research on models predicting acute kidney injury (AKI) [17] over 9 years has shown that while discriminatory performance (e.g., ranking outcomes) may remain stable, calibration drift (the misalignment of predicted probabilities with actual outcomes) can significantly undermine their clinical utility. Similar findings are reported in [18], where multiple regression and machine learning models exhibit calibration drift in the task of predicting 30-day hospital mortality. These emphasize the importance of performance monitoring to detect and correct degradation, ensuring AI models maintain their accuracy and relevance in dynamic healthcare environments.
@@
Even the most advanced AI models are not immune to unexpected performance degradation in real-world applications. The rise of large language models (LLMs) has marked a significant breakthrough in AI research, attracting global attention for their impressive capabilities. However, the increasing integration of LLMs into medicine and clinical workflows highlights the urgent need for robust AI performance monitoring. A recent study [19] assesses GPT-4’s performance on the American college of Radiology in-training examination and finds that when the questions were repeated several months later to assess for model drift, GPT-4 chose a different answer 25.5% of the time, indicating substantial temporal variability in LLM performance. A similar trend is observed in radiology question answering and interpretation tasks, where GPT-4 also demonstrates performance degradation over time [20]. A broader evaluation [21] highlights how GPT-4 and GPT-3.5 exhibit performance drift across various tasks, including code generation, math problem-solving, and sensitive question answering. One key observation is that GPT-4’s ability to follow user instructions degrades over time, contributing to the observed performance drops. These studies underscore the fact that even the most advanced AI systems can exhibit unexpected performance degradation in real-world scenarios across both clinical and non-clinical domains, reinforcing the necessity for continuous monitoring to ensure long-term reliability and safety.
@@
Figure 2: Overview of the medical AI monitoring framework, from performance degradation detection, analysis to correction.
@@
Our study is motivated by the growing use of AI in the medical and healthcare fields. As these systems are increasingly applied to real-world clinical settings, concerns about their long-term safety and reliability are becoming more important [22]. In particular, there is a growing awareness that monitoring AI after deployment deserves more attention [23, 24]. However, despite this need, there is still no comprehensive review that looks at the full picture of AI monitoring in medicine and healthcare, from detecting performance issues, to understanding their causes, to applying correction methods. To help address this gap, as shown in Fig. 2, our paper presents an integrated survey of existing techniques for detecting, analyzing, and correcting AI performance degradation in healthcare. By offering a clear and structured overview, we hope to support the development of AI systems that are not only effective at the start but also stay safe and reliable over time in real clinical environments.
@@ 2 Problem Formulation
Ensuring the long-term safety and reliability of ML models requires more than pre-deployment evaluation on training and test sets. It also demands robust post-deployment monitoring. Real-world environments often undergo changes in input data and usage context, which can cause performance degradation or unexpected failures if left unmanaged. We define AI performance monitoring as comprising two key components: data monitoring and model monitoring, both framed as distributional comparisons between datasets or model behaviors.
@@ 3.1.3 Concept Shift
Figure 3: Types of data shift in a disease classification model. The model classifies positive (disease, red) and negative (normal, green) cases. (a) Original distribution with a defined classification threshold. (b) Covariate shift: changes in input feature distribution (e.g., due to population differences). (c) Label shift: increase in disease prevalence. (d) Concept shift: changes in the relationship between features and labels.
@@ 3.1.4 Case Studies
The study in [27] explores the impact of data drift on machine learning models deployed in healthcare. Through simulations, it investigates covariate shift, concept shift, and major events (i.e., the COVID-19 pandemic) as sources of drift. The study highlights that data drift significantly impacts model performance, necessitating monitoring and retraining to maintain effectiveness.
@@ 3.2 Data Anomaly in Medical AI
In contrast to data shifts or drifts, which represent distribution-level changes, data anomalies refer to instance-level irregularities in input data that can cause sudden and severe model failure. These anomalies occur at the point level and are not necessarily associated with global distributional trends.
Data shift or drift is considered as a “soft failure” factor as models typically experience a gradual performance degradation due to slow changes in input distribution over time or across environments. In contrast, data anomalies are often sudden and severe, acting as “hard failure” triggers where even a single anomalous input can lead to unpredictable or harmful model behavior.
@@ 3.2.2 Definition and Types of Data Anomalies
Data anomalies encompass a wide variety of problematic inputs, including but not limited to:
  • Out-of-Distribution (OOD) Inputs: Samples that lie far from the training data manifold.
  • Highly Noisy Inputs: Corrupted or low-quality data that reduce signal-to-noise ratio.
  • Missing or Incomplete Data: Input with absent features, null values, or default placeholders.
  • Highly Biased Inputs: Data points that are unrepresentative or imbalanced, such as overrepresentation from a specific group, condition, or outlier case.
  • Adversarial Inputs: Carefully crafted samples designed to fool the model.
@@ 3.2.3 Impact on Model Performance
Data anomalies can cause sudden and serious model failures, unlike distribution shifts that usually lead to gradual performance drops (e.g., lower AUC or poor calibration). Even one bad input, like an extreme lab result or unreadable image, can lead to confident but wrong predictions, risking harmful clinical decisions.
While this survey focuses on shift-based failures that develop gradually and are often difficult to detect, sudden anomaly-induced errors are equally important. Robust AI monitoring should address both gradual distributional drift and abrupt data anomalies.
@@ 4.1 Calibration Drift
Calibration assesses how well predicted probabilities match actual outcomes [28]. In a well-calibrated model, a prediction of 0.8 corresponds to an 80% observed event rate. In healthcare, calibration is crucial for ensuring that predicted risks are both interpretable and actionable in clinical decision-making. However, when a model is deployed in new settings with different populations or disease prevalence, calibration can drift. This may lead to systematic overestimation or underestimation of risk even if the model’s ability to rank individuals by risks (i.e., discrimination or decision thresholds) remains unchanged. Such drifts can compromise the reliability of risk estimates and lead to suboptimal or inequitable clinical decisions.
@@ 4.2 Model Optimization Trade-Offs
Optimizing models for specific metrics-such as inference speed, computational efficiency, or adherence to safety constraints-can inadvertently degrade performance in other areas. For instance, the study by [19] suggests that GPT-4’s performance drift in radiology diagnosis tasks may be linked to optimization on other competing metrics, which potentially compromises diagnostic accuracy. These trade-offs can lead to unintended consequences in clinical AI models where accuracy and reliability are paramount.
A broader study [21] demonstrates GPT-4’s performance decline across multiple tasks over time, underscoring the inherent difficulty of maintaining consistent performance. It highlights how fine-tuning or system updates aimed at improving specific capabilities can cause unintended performance declines in other tasks.
@@ 4.3 Catastrophic Forgetting
To ensure adaptability in dynamic clinical environments, medical AI models are often updated with new patient data, shifting clinical practices, or evolving guidelines [29]. However, such update can unintentionally lead to catastrophic forgetting [30, 31], where the model loses performance on previously learned knowledge. Importantly, this does not only affects past cases but also future cases that share highly similar patterns with those older examples. In high-stakes settings like healthcare, catastrophic forgetting of previously learned disease subtypes, rare conditions, or specific population subgroups can create critical blind spots with potentially harmful consequences. Mitigating this risk requires careful update strategies, such as rehearsal-based methods, regularization techniques, or freezing critical model components [32].
@@ 4.5 Prompt Sensitivity
For models like LLMs, few-shot learning has demonstrated impressive performance in adapting to new tasks with minimal examples. However, studies [40] reveal that some LLMs and Vision-Language Models (VLMs) in medical domain can suffer performance degradation even after few-shot prompting, raising concerns about their robustness and real-world reliability. Additionally, prompt sensitivity, where variations in prompts lead to significantly different outputs, can result in inconsistencies and unpredictable behavior over time. Recent studies [41, 42, 43] further show that current LLMs often struggle to fully capture and manage this sensitivity. This limitation is particularly critical in high-stakes domains like healthcare, where reliability, reproducibility, and consistency are paramount.
@@ 5 Monitoring AI: Data Shift Detection
In medical AI systems, performance degradation can result from changes in input data (external) or the model itself (internal). Therefore, monitoring should cover both data inputs and model outputs. As shown in Fig. 4, this involves two key components: input monitoring (for data shift) and output monitoring (for model drift).
This section focuses on data shift detection. While not all data shifts cause performance drops, they are a leading contributor [44]. Detecting data shifts early can play a key role in tracking and maintaining model performance.
Figure 4: Illustration of AI performance monitoring: input (data) shift detection and output (model drift) detection.

### v2

# Keeping Medical AI Healthy and Trustworthy: A Review of Detection and Correction Methods for System Degradation

@@ Abstract
AI in Healthcare, AI Performance Degradation, AI Self-Correction, Data Shift, Model Drift, Clinical Decision Support, AI Monitoring, AI Reliability, AI Safety
@@ 1 Introduction
The world is seeing increasing applications of Artificial Intelligence (AI) in healthcare and medicine [1, 2, 3], transforming disease diagnosis, patient monitoring, outcome prediction, and treatment planning. AI technologies range from image analysis algorithms in radiology [4] and pathology [5] to predictive models that leverage electronic health records (EHR) for early disease detection [6]. By March 2025, the United States Food and Drug Administration (FDA) has listed 1,016 AI-enabled medical devices that have met the FDA’s applicable premarket requirements [7], and the number of AI/ML-based algorithms approved for clinical use by FDA continues to grow at a rapid rate [8]. As these AI systems continue to be adopted, ensuring their reliability and sustained performance becomes critically important. A recent study [9] analyzed 100 commercially available AI products in radiology and found that while AI adoption is accelerating, many systems lack rigorous scientific validation. Specifically, 64% of the products had no peer-reviewed evidence, and only a small fraction demonstrated clinical impact beyond diagnostic accuracy. The reliability of AI models in healthcare is not merely a technological concern but a crucial factor in ensuring patient safety and quality of care. Inaccurate or degraded AI performance can lead to severe consequences, including misdiagnosis, inappropriate treatment recommendations, and a loss of trust from healthcare providers and patients.
Figure 1: Illustration of performance degradation of medical AI models under temporal or spatial variations. The x-axis represents differences across time, or sites (proxy for data shift). Training data (left) include imaging, demographic, and EHR modalities, whereas deployment data (right) exhibit distributional changes (“shifted data”). Such shifts, regardless of their source, lead to reduced model accuracy and degraded performance, highlighting the importance of continuous AI monitoring.
@@
Another type of variation is within-environment variation, also known as temporal variation. In this scenario, an AI model is designed and deployed within a fixed environment, meaning the external setting remains constant. However, over time, the model still experiences performance degradation. This phenomenon can be vividly depicted as “AI aging” [13]. Please note this degradation arises from evolving variations in the underlying data distribution rather than time itself as a causal factor. Here, time serves as a proxy for data or model drift. Such temporal variations can occur abruptly, triggering a sharp drop in performance that can be classified as an anomaly event. Conversely, a more subtle but persistent issue arises when the AI model suffers from a gradual performance decline, which may go unnoticed without continuous monitoring.
Only recently has the research community and industry started recognizing the importance of monitoring AI performance degradation [14, 15, 16]. Chen et al.[17] demonstrate that clinical decision support systems trained on historical EHR data experience significant declines in predictive accuracy as the underlying clinical practice evolves, suggesting a natural decay in the utility of past clinical data. Their findings show that AI models relying on older datasets performed worse than those trained on more recent data, indicating an inherent susceptibility to performance degradation.
The study [18] empirically evaluates the long-term performance degradation of ML models predicting in-hospital mortality at the time of emergency admission, using a real-world dataset of 1.83 million patient discharge records from Maryland State Inpatient Database (2016–2018). By analyzing four top-performing ML models over 2.5 years, the authors track degradation across key metrics, including Area Under the Receiver Operator Characteristic Curve (AUROC), accuracy, precision, and recall. The findings reveal that while ML models can remain effective for over a year post-training, gradual performance decline necessitates strategic retraining. Additionally, the study highlights the limitations of 10-fold cross-validation in predicting long-term reliability and emphasizes the need for adaptive monitoring to maintain robustness in clinical AI applications.
@@
Another research on models predicting acute kidney injury (AKI) [20] over 9 years has shown that while discriminatory performance (e.g., ranking outcomes) may remain stable, calibration drift (the misalignment of predicted probabilities with actual outcomes) can significantly undermine their clinical utility. Similar findings are reported in [21], where multiple regression and machine learning models exhibit calibration drift in the task of predicting 30-day hospital mortality. These emphasize the importance of performance monitoring to detect and correct degradation, ensuring AI models maintain their performance in dynamic healthcare environments.
@@
Even the most advanced AI models are not immune to unexpected performance degradation in real-world applications. The rise of large language models (LLMs) has marked a significant breakthrough in AI research, attracting global attention for their impressive capabilities [22]. However, the increasing integration of LLMs into medicine and clinical workflows highlights the urgent need for robust AI performance monitoring. A recent study [23] assesses GPT-4’s performance on the American college of Radiology in-training examination and finds that when the questions were repeated several months later to assess for model drift, GPT-4 chose a different answer 25.5% of the time, indicating substantial temporal variability in LLM performance. A similar trend is observed in radiology question answering and interpretation tasks, where GPT-4 also demonstrates performance degradation over time [24]. A broader evaluation [25] highlights how GPT-4 and GPT-3.5 exhibit performance degradation across various tasks, including code generation, math problem-solving, and sensitive question answering. One key observation is that GPT-4’s ability to follow user instructions degrades over time, contributing to the observed performance drops. These studies underscore the fact that even the most advanced AI systems can exhibit unexpected performance degradation in real-world scenarios across both clinical and non-clinical domains, reinforcing the necessity for continuous monitoring to ensure long-term reliability and safety.
@@
Figure 2: Overview of the medical AI monitoring framework, from performance degradation detection, diagnosis to correction.
@@
Our study is motivated by the growing use of AI in the medical and healthcare fields. As these systems are increasingly applied to real-world clinical settings, concerns about their long-term safety and reliability are becoming more important [26]. In particular, there is a growing awareness that monitoring AI after deployment deserves more attention [27, 28]. However, despite this need, there is still no comprehensive review that looks at the full picture of AI monitoring in medicine and healthcare, from detecting performance issues, to understanding their causes, to applying correction methods.
To address this gap, our survey organizes existing methods for detecting, diagnosing, and correcting AI performance degradation in healthcare within a unified Detection-Diagnosis-Correction (DDC) cycle (Fig. 2). The DDC framework aligns closely with the FDA’s Predetermined Change Control Plan (PCCP) [29] for AI-enabled device software: “Detection” provides monitoring signals and predefined triggers outlined in the “Modification Protocol”, grounded in PCCP’s “Data Management Practices” and “Performance Evaluation” components; “Diagnosis” corresponds to “Impact Assessment” and “Traceability”, linking each proposed modification to its supporting protocol elements; “Correction” involves executing “Re-Training Practices”, verifying outcomes through performance evaluation, and implementing “Update Procedures”. In parallel, the DDC framework supports ISO/IEC 42001 [30] audit trails by reinforcing lifecycle documentation, decision transparency, and risk control. Collectively, this structure is designed to help medical AI systems remain safe, effective, and trustworthy throughout their deployment lifecycle.
@@ 2 Problem Formulation
Ensuring the long-term reliability and robustness of medical AI models requires more than pre-deployment evaluation on training and test sets. To avoid ambiguity, we distinguish between these related but distinct concepts:
  • Reliability refers to the ability of a medical AI system to consistently deliver accurate and trustworthy performance across time, populations, and clinical settings, ensuring predictable behavior throughout its lifecycle.
  • Robustness refers to the ability of a model to maintain stable performance under unexpected perturbations, such as noisy or corrupted inputs, and distributional shifts.
In this survey, we view robustness as a necessary component of reliability: a robust model is more likely to be reliable, but reliability additionally requires continuous monitoring, calibration, and correction over time. Building on this foundation, we define AI performance monitoring as comprising two key components: data monitoring and model monitoring, both framed as distributional comparisons between datasets or model behaviors. To avoid ambiguity, we distinguish these two components. Data monitoring focuses on identifying distributional changes in the input space (e.g., covariate shift), which may eventually lead to degraded model performance. In contrast, model monitoring targets performance decay or behavioral drift of the model itself. Importantly, model drift may stem from, but is not limited to, data drift. It can also arise from architectural or optimization-related factors. This separation ensures that our taxonomy captures both external changes in data and internal changes in model behavior.
@@ 2.2 Model Monitoring
2.3 Operational Parameters Selection for Drift Detection
Drift detection typically requires setting two key parameters: a threshold and a window size. The threshold (e.g., $\epsilon$ for data shift in (3) or $\delta_{m}$ for model drift in (5)) defines when the difference between current and reference data is large enough to be considered meaningful. In practice, even in post-deployment scenarios where labels are delayed or unavailable, it can be determined from prior experiments or adjusted to maintain the false-alarm rate within an acceptable range (e.g., 1-5%). Thresholds can also be refined adaptively once delayed ground-truth labels become available.
The window size (w) determines how much data is compared at each monitoring step. A smaller window increases sensitivity to short-term changes, whereas a larger window improves stability by reducing noise. In practice, rolling windows containing hundreds of samples or data from the past 1-3 months are commonly used, although the optimal choice ultimately depends on the specific system and task.
@@ 3.1.3 Concept Shift
Figure 3: Types of data shift in a disease classification model. The model classifies positive (disease, red) and negative (normal, green) cases using a fixed classifier (dashed line). a) Original data distribution. b) Covariate shift: changes in the input feature distribution (e.g., due to imaging acquisition differences) while the classifier remains fixed. c) Label shift: changes in class priors (e.g., increase in disease prevalence) under a fixed classifier. d) Concept shift: changes in the relationship between features and target output.
@@ 3.1.4 Case Studies
The study in [33] explores the impact of data shift on machine learning models deployed in healthcare. Through simulations, it investigates covariate shift, concept shift, and major events (i.e., the COVID-19 pandemic) as sources of data shift. The study highlights that data shift significantly impacts model performance, necessitating monitoring and retraining to maintain effectiveness.
@@ 3.2 Data Anomaly in Medical AI
In contrast to data shifts, which represent distribution-level changes, data anomalies refer to instance-level irregularities in input data that can cause sudden and severe model failure. These anomalies occur at the point level and are not necessarily associated with global distributional trends.
Data shift is considered as a “soft failure” factor as models typically experience a gradual performance degradation due to slow changes in input distribution over time. In contrast, data anomalies are often sudden and severe, acting as “hard failure” triggers where even a single anomalous input can lead to unpredictable or harmful model behavior.
@@ 3.2.2 Definition and Types of Data Anomalies
Data anomalies encompass a wide variety of problematic inputs, including but not limited to: 1) Highly Noisy Inputs: Corrupted or low-quality data that reduce signal-to-noise ratio; 2) Missing or Incomplete Data: Input with absent features, null values, or default placeholders; 3) Highly Biased Inputs: Data points that are unrepresentative or imbalanced, such as overrepresentation from a specific group, condition, or outlier case; 4) Adversarial Inputs: Carefully crafted samples designed to fool the model.
@@ 3.2.3 Impact on Model Performance
Data anomalies can cause sudden and serious model failures, unlike distribution shifts that usually lead to gradual performance drops (e.g., lower AUC). Even one bad input, like an extreme lab result or unreadable image, can lead to confident but wrong predictions, risking harmful clinical decisions.
While shift-based failures develop gradually and are often difficult to detect, sudden anomaly-induced errors are equally important. Robust AI monitoring should address both gradual distributional drift and abrupt data anomalies.
@@ 4.1 Calibration Drift
Calibration assesses how well predicted probabilities match actual outcomes [34]. In a well-calibrated model, a prediction of 0.8 corresponds to an 80% observed event rate. In healthcare, calibration is crucial for ensuring that predicted risks are both interpretable and actionable in clinical decision-making. However, when a model is deployed in new settings with different populations or disease prevalence, calibration can drift. This can distort risk estimates even when the model’s discrimination remains unchanged, undermining reliability and leading to suboptimal or inequitable clinical decisions.
@@ 4.2 Model Optimization Trade-Offs
Optimizing models for specific metrics, such as inference speed, computational efficiency, or adherence to safety constraints, can inadvertently degrade performance in other areas. For instance, the study by [23] suggests that GPT-4’s performance drift in radiology diagnosis tasks may be linked to optimization on other competing metrics, which potentially compromises diagnostic accuracy. These trade-offs can lead to unintended consequences in clinical AI models where accuracy and reliability are paramount.
A broader study [25] demonstrates GPT-4’s performance decline across multiple tasks over time, underscoring the inherent difficulty of maintaining consistent performance. It highlights that fine-tuning or system update aimed at improving specific capabilities can cause unintended performance declines in other tasks.
@@ 4.3 Catastrophic Forgetting
To ensure adaptability in dynamic clinical environments, medical AI models are often updated with new patient data, clinical practices, or evolving guidelines [35]. However, such update can unintentionally lead to catastrophic forgetting [36, 37], where the model loses performance on previously learned knowledge. Importantly, this does not only affects past cases but also future cases that share highly similar patterns with those older examples. In high-stakes settings like healthcare, catastrophic forgetting of previously learned disease subtypes, rare conditions, or specific population subgroups can create critical blind spots with potentially harmful consequences. Mitigating this risk requires careful update strategies, such as rehearsal-based methods, regularization techniques, or freezing critical model components [38].
@@ 4.5 Prompt Sensitivity
For models like LLMs, few-shot learning has demonstrated impressive performance in adapting to new tasks with minimal examples. However, studies [46] reveal that some LLMs in medical domain can suffer performance degradation after few-shot prompting, raising concerns about their robustness. Additionally, prompt sensitivity [47], where variations in prompts lead to significantly different outputs, can result in inconsistencies and unpredictable behavior over time. Recent studies [47, 48, 49] further show that current LLMs often struggle to fully capture and manage this sensitivity. This limitation is particularly critical in healthcare, where reliability, reproducibility, and consistency are paramount.
@@ 5 Monitoring AI: Data Shift Detection
In medical AI systems, performance degradation may arise from changes in input data (external) or the model itself (internal). Monitoring should therefore address both inputs and outputs, namely, input monitoring for data shift and output monitoring for model drift (Fig. 4). This section focuses on data shift detection, a primary source of performance decline [50, 51]. Early detection enables timely tracking and maintenance of model performance.
Figure 4: Illustration of AI performance monitoring: input (data) shift detection and output (model) drift detection.